OpenStudy (anonymous):

Medal.c: Choose the equation that represents the graph below. Graph of a line passing through points negative 3 comma 0 and 0 comma 3 y = x - 3 y = -x + 3 y = -x - 3 y = x + 3

OpenStudy (mayankdevnani):

\[\large \bf (-3,0) ~and~(0,3)\] You have to only find that equation that which satisfies both the points. For to satisfy these points,plug the values in all the equations and if they satisfy the condition(RIGHT HAND SIDE),then that option will be the right answer.

OpenStudy (mayankdevnani):

@kaylamarie081 can you solve it ?

OpenStudy (mayankdevnani):

understood ?? @kaylamarie081

OpenStudy (anonymous):

Umm choice B?

OpenStudy (mayankdevnani):

no !!

OpenStudy (mayankdevnani):

try it again. BETTER LUCK NEXT TIME !

OpenStudy (mayankdevnani):

answer is option D

OpenStudy (mayankdevnani):

EXPLANATION :- \[\large \bf y=x+3\] Plug (-3,0),we get \[\large \bf 0=-3+3=0\]

OpenStudy (mayankdevnani):

\[\large \bf y=x+3\] plug (0,3),we get \[\large \bf 3=0+3=3\]

OpenStudy (mayankdevnani):

option D satisfies all the given conditions. So,option D is correct answer. :)@kaylamarie081

OpenStudy (mayankdevnani):

@kaylamarie081

OpenStudy (anonymous):

Thank you uuuuuu soo mucho

OpenStudy (mayankdevnani):

welcome :)